A block of m1=20.0 kg is placed on a plane incline at angle

α=37.1° and is connected to a second hanging block with mass

m2 by a cord passing over a small, frictionless pulley. The

coefficient of kinetic friction between the plane and the block is

0.30.


1
Expert's answer
2022-03-15T16:49:30-0400

a)

"m_2=m_1(\sin\alpha+\mu\cos\alpha)=16.8~kg,"

b)

"m_2=m_1(\sin\alpha-\mu\cos\alpha)=7.3~kg,"

c)

"7.3~kg<m_2<16.8~kg."